### v3

# Absence of antisymmetric tensor fields: Clue from Starobinsky model of $f(R)$ gravity

Abstract
One of the surprising aspects of the present Universe, is the absence of any noticeable observable effects of higher-rank antisymmetric tensor fields in any natural phenomena. Here, we address the possible explanation of the absence of the higher rank antisymmetric tensor fields within the framework of $f(R)$ gravity. We explore the cosmological evolution of the scalar degrees of freedom associated with higher curvature term in a general $f(R)$ gravity model $f(R)=R+\alpha_{n}R^{n}$. We show that while different cosmological parameters mimic standard behaviour at different epochs for different forms of higher curvature gravity (i.e. different values of n ), only Starobinsky model (n = 2) gives a natural justification for the invisibility of the signatures of the massless modes of higher rank antisymmetric fields. In contrast, for other models ($n\neq 2$), despite their agreement with standard cosmology, the scalar degree of freedom induces an enhancement in the coupling of the antisymmetric fields and thereby contradicts the observation. The result does not change even with the inclusion of the Cosmological Constant. Thus, our result reveals that among different $f(R)$ models, Starobinsky model successfully explains the suppression of the massless modes of higher rank antisymmetric tensor fields leading to their invisibility in the present universe.
1 Introduction
We haven’t observed any overt indications of antisymmetric tensor fields having an impact on anything that occurs naturally, which is a rather startling aspect of our Universe. In this context, the second-rank antisymmetric tensor field, known as the Kalb-Ramond (KR) field [1], has been studied extensively. Such field appears naturally to cancel gauge anomaly in superstring theory[2, 3]. It has also been shown that the KR field has a natural geometric interpretation as space-time torsion in background geometry through an antisymmetric extension of the affine connection[4, 5, 6]. While it is simple to argue that the coupling of the KR field to matter should be $1/M_{p}$ in Einstein gravity, which is the same as the coupling of gravity with matter, based on dimensional considerations, there hasn’t been any experimental signature of the KR field in the observable Universe. As a result, it can be inferred that if such tensor fields exist, they must be greatly suppressed at the current scale of the Universe.
It was shown earlier that the presence of warped extra spatial dimension may lead to suppression of antisymmetric tensor field of various ranks[7, 8, 9, 10, 11]. While any extra-dimensional theory brings in a plethora of other predictions hitherto unobserved, within the domain of 3+ 1 dimensional Einstein gravity , it is hard to explain the reason of the disappearance of the massless KR field in our observable universe. As an alternative, in this work, we want to explore a dynamical mechanism of this suppression of the KR field as well as the other higher rank anti-symmetric fields in the light of higher curvature $f(R)$ theory. The presence of suitable higher curvature terms in the gravity action is permissible as long as they satisfy diffeomorphic invariance [12, 13, 14, 15, 16]. In this context $f(R)$ theories invoked great interest in the context of various cosmological predictions so far unexplained[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Higher order curvature terms in the gravity actions in such models are automatically suppressed by Planck scale which explains why at low energy only Einstein’s term dominates over the others. However, at high energy or at the epoch of early universe , the higher curvature term will have significant presence and in this context Starobinsky model [29] has been studied extensively in different context[30, 31, 32, 33, 34, 35, 36]. In this work, we explore whether the dynamical evolution in the presence of higher curvature gravity in 3+1 dimension may lead to a natural geometric suppression of the massless higher rank antisymmetric tensor fields.
In section 2, we present the standard formalism of transformation of an action in $f(R)$ gravity to Einstein gravity through a conformal transformation which manifests the higher curvature in terms of a scalar degree of freedom along with a potential. This transformation has been extensively discussed in the literature[12, 13, 17]. In section 3, we present the field equations of the transformed action in an isotropic and homogeneous universe. The conformal transformation induces coupling of scalar field with different matter. The coupled nonlinear equations are presented as a set of autonomous system in terms of a new set of dynamical variables. The form of $f(R)$ and the potentials generated from this modified gravity are discussed in section 4. Section 5 presents the solutions of the field equations in three different cases - A) $n=2$ or the Starobinsky model, B) $n\neq 2$ or generalised Starobinsky model and C) Both the models with Cosmological constant. In section 6 we highlight the finding of our work and finally, we discuss and conclude in section 7.
2 $f(R)$ Formalism and Conformal Transformation
For a rank $n$ antisymmetric tensor field $X_{a_{1}a_{2}...a_{n}}$, the field strength tensor of the massless mode can be written as $Y_{a_{1}a_{2}...a_{n+1}}=\partial_{[a_{n+1}}X_{a_{1}a_{2}...a_{n}]}$. In 3+1 dimensional spacetime, the maximum rank of an antisymmtric tensor field can be 3 as the field strength vanishes beyond that. The action of higher curvature gravity in 3+1 dimension along with other fields can be written as
$\begin{split}S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\bigg{(}\frac{F(R)}{2\kappa^{2}}+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\bigg{)}\end{split}$ (2.1)
where the last 4 terms are the kinetic lagrangian of spin $\frac{1}{2}$ fermion fields $\psi$, U(1) gauge field $A_{\mu}$, Kalb Ramond field $B_{\mu\nu}$ and rank 3 antisymmetric tensor field $X_{\mu\nu\rho}$ respectively. $H_{\mu\nu\rho}$ and $T_{\mu\nu\rho\delta}$ are the field strength tensor of the Kalb Ramond field and rank 3 antisymmetric tensor field respectively defined as $H_{\mu\nu\rho}=\partial_{[\mu}B_{\nu\rho]}$ and $T_{\mu\nu\rho\delta}=\partial_{[\mu}X_{\nu\rho\delta]}$.
It is standard to represent [12, 13] the action (2.1) in the following way with an introduction of an auxiliary field
$S=$ $\int d^{4}x\sqrt{-g}\Bigg{(}\frac{1}{2\kappa^{2}}\phi R-V(\phi)+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\Bigg{)}$ (2.2)
where
$\phi\equiv f^{\prime}(R)~{}\text{and}~{}V(\phi)=\frac{1}{2\kappa^{2}}[{\phi~{}R-f(R)}]$ (2.3)
Now, following a conformal transformation of the form $\tilde{g}_{\mu\nu}(x)=\Omega^{2}{g}_{\mu\nu}(x)$, where $\Omega^{2}=f^{\prime}=\phi$ one can rewrite the action (2.2) in the Einstein frame as
$S_{E}=$ $\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})+e^{\sqrt{\frac{2}{3}}k\tilde{\phi}}\Psi^{+}\tilde{\gamma}^{0}i\tilde{\gamma}^{\mu}D_{\mu}{\Psi}-\frac{1}{4}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}{F}_{\mu\nu}{F}_{\alpha\beta}$ (2.4)
$-\frac{1}{4}e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}\widetilde{g}^{\rho\delta}{H}_{\mu\nu\rho}{H}_{\alpha\beta\delta}-\frac{1}{4}e^{-2\sqrt{\frac{2}{3}}k\tilde{\phi}}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}\widetilde{g}^{\rho\delta}\widetilde{g}^{\sigma\lambda}{T}_{\mu\nu\rho\sigma}{T}_{\alpha\beta\delta\lambda}\bigg{)}$
where the scalar field $\tilde{\phi}$ is defined as $\kappa\tilde{\phi}\equiv\sqrt{\frac{3}{2}}\ln\phi$ and the metric in Einstein frame is represented by an overhead $\sim$. This choice of conformal transformation is consistent with $f^{\prime}>0$.
It should be noted that in the last two terms, which are the field strength tensors of the massless mode of the rank 2 and 3 antisymmetric tensor fields, an exponential factor of $\tilde{\phi}$ is present. So, the possible signatures of these field crucially depends on the nature of the scalar field $\tilde{\phi}$ in these exponential factors
(i) If the scalar field $\tilde{\phi}$ is positive then the exponential factor provides an additional suppression to these terms making the signature of the anti symmetric fields invisible.
(ii) Otherwise if $\tilde{\phi}$ is negative, the exponential factor will enhance the overall coupling, thereby making the signature significant.
Hence we study the evolution of the scalar field $\tilde{\phi}$ to understand the nature of the field which in turn will determine whether the imprints of the higher rank antisymmtric tensor fields will be visible or not.
3 Field Equations in Einstein Frame
Our goal is to explore how the scalar field evolves cosmologically in the Einstein frame in a background FRW space-time in order to understand whether the coupling of the higher rank antisymmetric fields are suppressed or enhanced as the universe evolves. And to do so, we further rewrite the entire action into two parts - one part consists of the scalar field ($\tilde{\phi}$) in Einstein gravity and the other part contains all other matter fields, namely, fermions, gauge fields, and higher rank antisymmetric fields which are coupled to the scalar field $\tilde{\phi}$ through the metric $g_{\mu\nu}$. Note that we do not consider any back reaction of these fields on the scalar field.
$S_{E}$ $=\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})\bigg{)}$ (3.1)
$+\int d^{4}x~{}{\cal L}_{M}(e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}(x)}\tilde{g}_{\mu\nu},{\psi},{A}_{\mu},{B}_{\mu\nu},{X}_{\mu\nu\rho})$
${{\cal L}_{\tilde{\phi}}}=\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})$ is the Lagrangian density of the scalar field ${\tilde{\phi}}$ and ${{\cal L}_{M}}$ is the Lagrangian density of matter fields. The energy-momentum tensor of the scalar field and matter, respectively is defined as,
${\tilde{T}^{\tilde{\phi}}}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial(\sqrt{-\tilde{g}}{{\cal L}_{\tilde{\phi}}})}{\partial{\tilde{g}}^{\mu\nu}}~{}~{}\text{and}~{}~{}{\tilde{T}}^{M}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{g}}^{\mu\nu}}$ (3.2)
Variation of the action (3.1) with respect to the field $\tilde{\phi}$ gives the dynamical equation for the scalar field
$\square{\tilde{\phi}}-\frac{\partial V}{\partial{\tilde{\phi}}}+\frac{1}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=0$ (3.3)
where ${\cal{L}}_{M}\equiv{\cal{L}}_{M}(e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}(x)}\tilde{g}_{\mu\nu})$. Hence
$\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=\frac{\partial{\cal L}_{M}}{\partial g^{\mu\nu}}\frac{\partial g^{\mu\nu}}{\partial{\tilde{\phi}}}={\sqrt{-\tilde{g}}}\kappa Q{\tilde{T}}$ (3.4)
where $Q=-\frac{1}{\sqrt{6}}$ is the strength of the coupling between the field and matter, which is constant for any form of $f(R)$.
The matter in (3.2) is considered in the form of perfect fluid $\tilde{T}^{\mu(M)}_{\nu}=\text{diag}(-\tilde{\rho}_{M},\tilde{P}_{M},\tilde{P}_{M},\tilde{P}_{M})$. $\tilde{\rho}_{M}$ and $\tilde{P}_{M}$ are the energy density and pressure of matter. We consider the observable universe to consist of radiation and matter (baryonic and CDM) i.e., $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}$.
In a flat FRW background the field equations are given by
$3\tilde{H}^{2}=\kappa^{2}[\frac{1}{2}(\frac{d\tilde{\phi}}{d\tilde{t}})^{2}+V(\tilde{\phi})+\tilde{\rho}_{M}]$ (3.5)
$\frac{d^{2}\tilde{\phi}}{d\tilde{t}^{2}}+3\tilde{H}\frac{d\tilde{\phi}}{d\tilde{t}}+\frac{\partial V(\tilde{\phi})}{\partial{\tilde{\phi}}}=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})$ (3.6)
$\frac{d\tilde{\rho}_{M}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{M}+\tilde{P}_{M})=\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.7)
where $\tilde{H}\equiv\frac{1}{\tilde{a}}\frac{d\tilde{a}}{\tilde{d}t}$. One can rewrite equation (3.6) as
$\frac{d\tilde{\rho}_{\phi}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.8)
where $\tilde{\rho}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})+V(\tilde{\phi}))$ is the energy density of the scalar field and $\tilde{P}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})-V(\tilde{\phi}))$ is the pressure of the scalar field Using the standard thermodynamic relation $\tilde{P}_{M}=\omega\tilde{\rho}_{M}$, one can easily find a relation between $\tilde{\rho}_{M}$ and $\tilde{\phi}$ from equation (3.7)
$\tilde{\rho}_{M}=\tilde{\rho}_{M0}\tilde{a}^{-3(1+\omega)}\exp({\frac{\kappa}{\sqrt{6}}(1-3\omega)(\tilde{\phi}-\tilde{\phi}_{0})})$ (3.9)
So the set of dynamical equations to be handled are
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}\bigg{]}$ (3.10)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m})\tilde{\phi}^{\prime}$ (3.11)
$\tilde{\rho}_{m}^{\prime}+3\tilde{\rho}_{m}=\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{m}\tilde{\phi}^{\prime}$ (3.12)
$\tilde{\rho}_{r}^{\prime}+4\tilde{\rho}_{r}=0$ (3.13)
where prime denotes the differentiation with respect to $N(=\ln{\tilde{a}})$. Note that since the radiation field is traceless the scalar field does not couple to radiation (3.4). In order to solve these coupled nonlinear inhomogeneous equations, we introduce a set of dynamical variables [37, 38, 39]
$x=\frac{\kappa\tilde{\phi}\prime}{\sqrt{6}},y=\frac{\kappa\sqrt{V}}{\sqrt{3}\tilde{H}},\lambda=-\frac{V_{,\tilde{\phi}}}{kV},~{}m=\frac{\kappa^{2}\tilde{\rho}_{m}}{3\tilde{H}^{2}},~{}r=\frac{\kappa^{2}\tilde{\rho}_{r}}{3\tilde{H}^{2}}$ (3.14)
and rewrite the equations (3.10-3.13) in terms of these new variables forming a set of autonomous systems of first-order differential equations
[…]
where, $\Gamma=\frac{VV_{\tilde{\phi}\tilde{\phi}}}{V_{\tilde{\phi}}^{2}}$. In terms of these new variables, equation (3.10) takes the form $1=x^{2}+y^{2}+m+r$.
4 General class of $f(R)$ : $f(R)=R+\alpha_{n}R^{n}$
In this work, we consider a very general class of modified gravity $f(R)=R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has a suitable dimension. The Starobinsky model of modified gravity corresponds to $n=2$ i.e., $f(R)=R+\alpha_{2}R^{2}$, where $\alpha_{2}\sim\frac{1}{M^{2}}$. With this general form of $f(R)$, the corresponding potential is
$V(\tilde{\phi})={\cal A}e^{-2\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\bigg{(}e^{\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}-1\bigg{)}^{\frac{n}{n-1}}$ (4.1)
where ${\cal A}=\frac{(n-1)\alpha_{n}}{2\kappa^{2}}\big{[}\frac{1}{n\alpha_{n}}\big{]}^{\frac{n}{n-1}}$. The potential becomes zero at $\tilde{\phi}=0$, irrespective of $n$. For $n=2$, the potential is constant when $\tilde{\phi}$ is positive. For negative $\tilde{\phi}$ it increases exponentially. For $n\neq 2$, the potential increases sharply for negative $\tilde{\phi}$ and goes to 00 for positive $\tilde{\phi}$. In fig (1) we present the form of potentials for various values of $n$ with ${\cal A}=1$.
Figure 1: Form of potentials for different $n$
For the general potential $\Gamma$ takes a nontrivial form
$\Gamma=\frac{\frac{n}{(n-1)^{2}}-\frac{3n}{n-1}\left[1-e^{-\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\right]+4\left[1-e^{-\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\right]^{2}}{\left[\frac{n}{n-1}-2(1-e^{-\sqrt{\frac{2}{3}}\kappa\tilde{\phi}})\right]^{2}}$
We solve these equations numerically taking different integral values of $n$ ranging from -4 to 4 with suitable initial conditions. Corresponding to different values of $n$, the expression for $\Gamma$ changes and hence the set of equations.
5 Different cases of $n$
5.1 Case I: $n=2$, Starobinsky model
For $n=2$, $f(R)$ takes the form $R+\alpha_{2}R^{2}$, which is the famous Starobinsky model. Substituting the value of $n$, the $\frac{d\lambda}{dN}$ equation becomes
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{2}}\lambda^{2}x\left[1-\frac{1}{\lambda}\sqrt{\frac{8}{3}}\right]$ (5.1)
Note that the other equations in (3.15) remain the same. We solve the above equations numerically, with suitable initial conditions such that radiation-matter equality happens around redshift 3300 and matter in the present universe constitutes approximately 30% of the energy density. In Fig (2), we present different density parameters, scalar field $\tilde{\phi}$ and the equation of state $\omega_{\tilde{\phi}}=\frac{\tilde{P}_{\tilde{\phi}}}{\tilde{\rho}_{\tilde{\phi}}}$ with respect to $N(=\ln{\tilde{a}})$ The density parameters of different components mimic the behaviours in standard model of cosmology. From the plot it is evident that the energy density of $\tilde{\phi}$ field never dominates the evolution until recently.
Figure 2: Evolution of density parameters, scalar field $\tilde{\phi}$ and $\omega_{\tilde{\phi}}$ in Starobinsky model
$\tilde{\phi}$ remains positive along the entire evolution and does not vary much. $\omega_{\tilde{\phi}}$ remains constant for a long period and starts varying late, around $z\sim 7$ and finally approaches $-1$ at present. So, in the late time the field behaves as dark energy. This behaviour of $\omega_{\tilde{\phi}}$ resembles the equation of state of the thawing quintessence field.
We further investigate the potential and find that the scalar field rolls down the potential $V(\tilde{\phi})$ but does not reach the minima. This is due to the coupling between the scalar field and matter (see equations (3.11) and (3.12)). Plotting both the terms separately, we realised that the potential arising due to coupling with matter $V_{m}$ is far more dominant than the potential arising from modified gravity $V(\tilde{\phi})$. Hence the scalar field did not reach the minima. In Fig (3), we present the three different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$.
Figure 3: Different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$ in Starobinsky model
5.2 Case 2: $n\neq 2$, $f(R)=R+\alpha_{n}R^{n}$
Similarly, we have studied the other cases where $n$ takes positive values $3,4$ as well as negative values $-1,-2,-3,-4$ in the modified gravity form $f(R)=R+\alpha_{n}R^{n}$ and suitably changed the set of dynamical equations. We have already noticed that for all these cases, the potential form remains the same irrespective of different values of $n$. This hints that for all other cases $(n\neq 2)$ the results are expected to be same. Hence we present here only one case ($n=-2$) as an example. For $n=-2$, the $\frac{d\lambda}{dN}$ equation becomes
$\frac{d\lambda}{dN}=\sqrt{\frac{3}{8}}x(4\sqrt{8/3}-6\lambda)(\sqrt{8/3}-\lambda)$ (5.2)
along with the other equations in equation(3.15). Following similar methods as mentioned above, we solve the equations numerically, with suitable initial conditions. We did a similar analysis of the cosmological parameters as we have done in the case of $n=2$. Though the form of the potential of $V(\phi)$ in this case is different from the Starobinsky model,interestingly, we find that the behavior of the density parameters and the equation of state almost exactly matches the Starobinsky model over the entire period of evolution (Fig 4). This is because here also the potential $V_{m}$ arising due to matter coupling dominates, hence the effective potential $V_{eff}$ almost similarly matches the case of $n=2$ (Fig 5). The only difference is that the scalar field is negative for the entire period as it rolls down the potential. For all other values of n, ($\neq 2$), we get the same results.
Figure 4: Evolution of different cosmological parameters for $n=-2$ case
Figure 5: Different potentials for $n=-2$ case
5.3 $f(R)=R+\alpha_{n}R^{n}$ models with Cosmological Constant
We have also carried out the entire study including the Cosmological Constant $\Lambda$ along with the other matter components in the observable universe i.e, $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}+\tilde{\rho}_{\Lambda}$. The scalar field $\tilde{\phi}$ also couples with $\Lambda$ also, apart from matter. So, equations (3.10) and (3.11) changes to include $\Lambda$, while (3.12) and (3.13) remain unchanged. Also, we have one more coupled equation for $\Lambda$.
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}+\tilde{\rho}_{\Lambda}\bigg{]}$ (5.3)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m}+4\tilde{\rho_{\Lambda}})\tilde{\phi}^{\prime}$ (5.4)
$\tilde{\rho}_{\Lambda}^{\prime}=4\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{\Lambda}\tilde{\phi}^{\prime}$ (5.5)
To proceed with the analysis we define one more dynamical variable $l=\frac{\kappa^{2}\tilde{\rho}_{\Lambda}}{3\tilde{H}^{2}}$. In terms of all the dynamical variables, the new set of equations to be dealt with looks like
$\frac{dx}{dN}$ $=\frac{x}{2}(3x^{2}+3-3y^{2}+r-3l)-3x+\sqrt{\frac{3}{2}}\lambda y^{2}$ (5.6)
$-\frac{1}{2}(1-x^{2}-y^{2}-r-l)-2l$
$\frac{dy}{dN}$ $=-y[\sqrt{\frac{3}{2}}x\lambda+\frac{1}{2}(3y^{2}-3-3x^{2}-r+3l)]$
$\frac{d\lambda}{dN}$ $=\sqrt{6}x\lambda^{2}[1-\Gamma]$
$\frac{dr}{dN}$ $=-r[4+(-3x^{2}+3y^{2}-r+3l-3)]$
$\frac{dl}{dN}$ $=l[4x-(-3x^{2}+3y^{2}-r+3l-3)]$
We perform a similar numerical analysis with the new set of equations for the potential given by (4.1) with different $n$ as has been done without $\Lambda$. Surprisingly, we found that inclusion of the Cosmological constant does not effect the evolution of different components. And like the earlier scenario, the coupling term with matter $V_{m}$ is far more dominant than the potential $V(\tilde{\phi})$ arising out of modified gravity as well as the coupling with the cosmological constant $V_{l}$. Hence driven mostly by the matter coupling term $V_{m}$, the evolution of different components remain similar. The value of the scalar field for different $n$ also matches our earlier findings. For $n=2$, scalar field remains positive throughout, while for other models $(n\neq 2)$ it remains negative even after inclusion of $\Lambda$. Here we present only one case $n=2$ as an example.
Figure 6: Evolution of different cosmological parameters and potentials including $\Lambda$ for $n=2$
6 Results
We have considered a very general class of modified gravity of the form $f(R)=R+\alpha_{n}R^{n}$ gravity, which is equivalent to Einstein gravity along with a scalar degree of freedom $\tilde{\phi}$ following the potential given by (4.1), when conformally transformed in Eintein frame. Depending on the behaviour of the potential we divided our cases in two categories $n=2$ (Case I) and $n\neq 2$ (Case 2). We investigated the evolution of different components like scalar field, radiation and matter in a background isotropic and homogeneous universe represented by FRW metric. After the whole analysis we found that in case $n=2$, we found that $\tilde{\phi}$ remains positive along the entire evolution and does not vary much. The matter - scalar field coupling does not allow the scalar field to reach the minima at zero. Such values of $\tilde{\phi}$ for the last two terms in action (2.4) provide an additional exponential suppression. This makes the signature of the antisymmetric fields (both rank 2 and 3) heavily suppressed thereby failing to leave some imprints in the observable universe. The investigations for the case of $n\neq 2$ reveal $\tilde{\phi}$ to be negative all along (again matter-$\tilde{\phi}$ coupling not allowing $\tilde{\phi}$ to reach zero). Such values of $\tilde{\phi}$ would imply enhancement of the coupling for the last two terms in action (2.4) by an exponential factor. This indicates that signatures of the antisymmetric fields of rank 2 and 3 should be present, which is contrary to the observation. For all other values of $n\neq 2$, we get the same results. We also performed the analysis adding Cosmological constant $\Lambda$ along with the earlier components in the FRW background. Even though the equations change significantly, the nature of the scalar field for different $n$ matches our earlier findings. So with or without $\Lambda$, the antisymmetric fields of rank 2 and 3 gets heavily suppressed in case of Starobinsky model $(n=2)$, justifying the invisibility of these fields, while their couplings gets enhanced for other cases $(n\neq 2)$, implying presence of such fields, contrary to the observation.
7 Discussion and Conclusion
In our current universe, we don’t see any signs of higher rank antisymmetric tensor fields having an impact on natural events. We explain the suppression of these fields in the light of the higher curvature $f(R)$ theory. It has been extensively discussed in literature that modified $f(R)$ gravity when expressed in the Einstein frame, is equivalent to Einstein gravity along with a scalar degree of freedom $(\tilde{\phi})$ following a potential, governed by the form of $f(R)$. Thus, in the Einstein frame, $\tilde{\phi}$ couples with all the other fields through the conformal transformation of the metric. In this work, we first studied how different components like scalar field, radiation and matter evolve in a background isotropic and homogeneous universe represented by FRW metric. Further, we explore how the nature of the scalar field influences the coupling of the massless modes of higher rank antisymmetric fields present in the action (2.4). The coupling, in turn, is responsible for the observation of the possible imprints of massless modes of these antisymmetric fields.
For this study, we have considered the form of $f(R)$ to be $R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has suitable dimensions. As shown in fig(1), the shape of the potential $V(\tilde{\phi})$ for this form is same for all values of $n$, apart from $n=2$. For $n\neq 2$, the potential is zero for positive $\tilde{\phi}$ and increases exponentially for $\tilde{\phi}<0$. The $n=2$ case, Starobinsky model, is different from the rest, as the potential increases and soon reaches a constant for positive $\tilde{\phi}$ but for negative $\tilde{\phi}$, the behaviour is the same as other cases. The field equations with different kinds of matter and scalar field in the homogeneous and isotropic background are coupled nonlinear differential equations. We rewrite the equations as a set of autonomous system of first order differential equations to tackle the complexity. The autonomous system is solved numerically, with suitable initial conditions such that, following standard model of cosmology, the matter radiation equality occurs around redshift $z\sim 3300$ and the matter content is $30\%$ of the total in the present universe.
Though we studied different cases for $-4\geq n\geq 4$, we broadly divide them into two categories $n=2$ and $n\neq 2$, depending on the behaviour of the potential. We found the evolution of all components mimic their behaviour in standard model of cosmology and the scalar field $\tilde{\phi}$ is subdominant throughout the evolution of the universe until recently. Further investigation of the equation of state $\omega_{\tilde{\phi}}$ reveals in the late time $\tilde{\phi}$ behaves as a source of dark energy, which resemble the behaviour of thawing quintessence model. Surprisingly we obtain similar behaviours for all of these parameters for both the categories $(n=2~{}{\text{and}}~{}n\neq 2)$, even though the potentials are different. The reason for such similarity is that the dynamics is dominated by the coupling between matter and $\tilde{\phi}$, which gives rise to an effective potential. The effective potential exhibit similar nature for different values of $n$ and the variation of potential $V(\tilde{\phi})$ does not have any significant impact on the evolution of the parameters.
Most interestingly, for different values of $n$, even though the different matter components, namely matter, radiation and scalar field evolve in the similar fashion with time, $\tilde{\phi}$ is positive only for $n=2$ and negative for $n\neq 2$, throughout. The value of $\tilde{\phi}$ depends on the slope of the potential. For $n=2$, $\tilde{\phi}$ remains positive as it rolls towards the minima ($\tilde{\phi}=0$). Where as, for $n\neq 2$, $\tilde{\phi}$ can roll only over negative value, as on the positive side the potential is zero. However, for both the cases $\tilde{\phi}$ can never reach zero due to coupling with matter. Hence even though both the categories agree with standard model of cosmology, only for Starobinsky model $(n=2)$, the last two terms of the action (2.4) get an extra exponential suppression. For $n\neq 2$, we obtain the enhancement of the coupling term instead of the suppression. As there has been no experimental evidence of the footprint of antisymmetric fields on the present Universe, so $n=2$ i.e., the Starobinsky model is the only model in a general class of modified gravity where $f(R)=R+\alpha_{n}R^{n}$, to support the suppression of such fields. We have also added cosmological constant $\Lambda$ as another component. Even though $\Lambda$ also couples with $\tilde{\phi}$, the effect is very less with respect to matter-$\tilde{\phi}$ coupling which remains the dominant term in the whole analysis. Hence, the dynamics and inference regarding the suppression of the antisymmetric fields remain unchanged. So with or without $\Lambda$ the Starobinsky model is the only model to justify the absence of signature of the massless modes of antisymmetric fields in a general class of modified gravity model.
8 Acknowledgement
Sonej Alam acknowledges the CSIR fellowship provided by the Govt. of India under the CSIR-JRF scheme (file no. 09/0466(12904)/2021). The authors thank Wali Hossain and Anjan A. Sen for their invaluable insights and suggestions.

### v4

@@
 Thus, our result reveals that among different $f(R)$ models, Starobinsky model successfully explains the suppression of the massless modes of higher rank antisymmetric tensor fields leading to their invisibility in the present universe.
+keywords:
+f(R) Theory, Einstein Frame, Dynamical Dark Energy, Torsion
+journal: Physics of the Dark Universe
+[…]
+[…]
 1 Introduction
@@
 However, at high energy or at the epoch of early universe , the higher curvature term will have significant presence and in this context Starobinsky model [29] has been studied extensively in different context[30, 31, 32, 33, 34, 35, 36].
-In this work, we explore whether the dynamical evolution in the presence of higher curvature gravity in 3+1 dimension may lead to a natural geometric suppression of the massless higher rank antisymmetric tensor fields.
+In this work, we explore whether the dynamical evolution of the components in the presence of higher curvature gravity in 3+1 dimension may lead to a natural geometric suppression of the massless higher rank antisymmetric tensor fields.
 In section 2, we present the standard formalism of transformation of an action in $f(R)$ gravity to Einstein gravity through a conformal transformation which manifests the higher curvature in terms of a scalar degree of freedom along with a potential.
@@
 In section 3, we present the field equations of the transformed action in an isotropic and homogeneous universe.
-The conformal transformation induces coupling of scalar field with different matter.
+The conformal transformation induces coupling of scalar field with different component.
 The coupled nonlinear equations are presented as a set of autonomous system in terms of a new set of dynamical variables.
@@
 In 3+1 dimensional spacetime, the maximum rank of an antisymmtric tensor field can be 3 as the field strength vanishes beyond that.
-The action of higher curvature gravity in 3+1 dimension along with other fields can be written as
-$\begin{split}S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\bigg{(}\frac{F(R)}{2\kappa^{2}}+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\bigg{)}\end{split}$ (2.1)
+In an early universe post reheating epoch (around the time of nucleosynthesis), the effective action of higher curvature gravity in 3+1 dimension along with other fields can be written as
+$\begin{split}S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\bigg{(}\frac{F(R)}{2\kappa^{2}}+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\bigg{)}\end{split}$ (1)
 where the last 4 terms are the kinetic lagrangian of spin $\frac{1}{2}$ fermion fields $\psi$, U(1) gauge field $A_{\mu}$, Kalb Ramond field $B_{\mu\nu}$ and rank 3 antisymmetric tensor field $X_{\mu\nu\rho}$ respectively.
 $H_{\mu\nu\rho}$ and $T_{\mu\nu\rho\delta}$ are the field strength tensor of the Kalb Ramond field and rank 3 antisymmetric tensor field respectively defined as $H_{\mu\nu\rho}=\partial_{[\mu}B_{\nu\rho]}$ and $T_{\mu\nu\rho\delta}=\partial_{[\mu}X_{\nu\rho\delta]}$.
-It is standard to represent [12, 13] the action (2.1) in the following way with an introduction of an auxiliary field
-$S=$ $\int d^{4}x\sqrt{-g}\Bigg{(}\frac{1}{2\kappa^{2}}\phi R-V(\phi)+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\Bigg{)}$ (2.2)
+It is standard to represent [12, 13] the action (1) in the following way with an introduction of an auxiliary field
+$S=$ $\int d^{4}x\sqrt{-g}\Bigg{(}\frac{1}{2\kappa^{2}}\phi R-V(\phi)+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\Bigg{)}$ (2)
 where
-$\phi\equiv f^{\prime}(R)~{}\text{and}~{}V(\phi)=\frac{1}{2\kappa^{2}}[{\phi~{}R-f(R)}]$ (2.3)
-Now, following a conformal transformation of the form $\tilde{g}_{\mu\nu}(x)=\Omega^{2}{g}_{\mu\nu}(x)$, where $\Omega^{2}=f^{\prime}=\phi$ one can rewrite the action (2.2) in the Einstein frame as
-$S_{E}=$ $\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})+e^{\sqrt{\frac{2}{3}}k\tilde{\phi}}\Psi^{+}\tilde{\gamma}^{0}i\tilde{\gamma}^{\mu}D_{\mu}{\Psi}-\frac{1}{4}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}{F}_{\mu\nu}{F}_{\alpha\beta}$ (2.4)
+$\phi\equiv f^{\prime}(R)~{}\text{and}~{}V(\phi)=\frac{1}{2\kappa^{2}}[{\phi~{}R-f(R)}]$ (3)
+Now, following a conformal transformation of the form $\tilde{g}_{\mu\nu}(x)=\Omega^{2}{g}_{\mu\nu}(x)$, where $\Omega^{2}=f^{\prime}=\phi$ one can rewrite the action (2) in the Einstein frame as
+$S_{E}=$ $\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})+e^{\sqrt{\frac{2}{3}}k\tilde{\phi}}\Psi^{+}\tilde{\gamma}^{0}i\tilde{\gamma}^{\mu}D_{\mu}{\Psi}-\frac{1}{4}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}{F}_{\mu\nu}{F}_{\alpha\beta}$ (4)
 $-\frac{1}{4}e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}\widetilde{g}^{\rho\delta}{H}_{\mu\nu\rho}{H}_{\alpha\beta\delta}-\frac{1}{4}e^{-2\sqrt{\frac{2}{3}}k\tilde{\phi}}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}\widetilde{g}^{\rho\delta}\widetilde{g}^{\sigma\lambda}{T}_{\mu\nu\rho\sigma}{T}_{\alpha\beta\delta\lambda}\bigg{)}$
@@
 Note that we do not consider any back reaction of these fields on the scalar field.
-$S_{E}$ $=\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})\bigg{)}$ (3.1)
+$S_{E}$ $=\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})\bigg{)}$ (5)
 $+\int d^{4}x~{}{\cal L}_{M}(e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}(x)}\tilde{g}_{\mu\nu},{\psi},{A}_{\mu},{B}_{\mu\nu},{X}_{\mu\nu\rho})$
@@
 The energy-momentum tensor of the scalar field and matter, respectively is defined as,
-${\tilde{T}^{\tilde{\phi}}}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial(\sqrt{-\tilde{g}}{{\cal L}_{\tilde{\phi}}})}{\partial{\tilde{g}}^{\mu\nu}}~{}~{}\text{and}~{}~{}{\tilde{T}}^{M}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{g}}^{\mu\nu}}$ (3.2)
-Variation of the action (3.1) with respect to the field $\tilde{\phi}$ gives the dynamical equation for the scalar field
-$\square{\tilde{\phi}}-\frac{\partial V}{\partial{\tilde{\phi}}}+\frac{1}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=0$ (3.3)
+${\tilde{T}^{\tilde{\phi}}}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial(\sqrt{-\tilde{g}}{{\cal L}_{\tilde{\phi}}})}{\partial{\tilde{g}}^{\mu\nu}}~{}~{}\text{and}~{}~{}{\tilde{T}}^{M}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{g}}^{\mu\nu}}$ (6)
+Variation of the action (5) with respect to the field $\tilde{\phi}$ gives the dynamical equation for the scalar field
+$\square{\tilde{\phi}}-\frac{\partial V}{\partial{\tilde{\phi}}}+\frac{1}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=0$ (7)
 where ${\cal{L}}_{M}\equiv{\cal{L}}_{M}(e^{-\sqrt{\frac{2}{3}}k\tilde{\phi}(x)}\tilde{g}_{\mu\nu})$. Hence
-$\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=\frac{\partial{\cal L}_{M}}{\partial g^{\mu\nu}}\frac{\partial g^{\mu\nu}}{\partial{\tilde{\phi}}}={\sqrt{-\tilde{g}}}\kappa Q{\tilde{T}}$ (3.4)
+$\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=\frac{\partial{\cal L}_{M}}{\partial g^{\mu\nu}}\frac{\partial g^{\mu\nu}}{\partial{\tilde{\phi}}}={\sqrt{-\tilde{g}}}\kappa Q{\tilde{T}}$ (8)
 where $Q=-\frac{1}{\sqrt{6}}$ is the strength of the coupling between the field and matter, which is constant for any form of $f(R)$.
-The matter in (3.2) is considered in the form of perfect fluid $\tilde{T}^{\mu(M)}_{\nu}=\text{diag}(-\tilde{\rho}_{M},\tilde{P}_{M},\tilde{P}_{M},\tilde{P}_{M})$.
+The matter in (6) is considered in the form of perfect fluid $\tilde{T}^{\mu(M)}_{\nu}=\text{diag}(-\tilde{\rho}_{M},\tilde{P}_{M},\tilde{P}_{M},\tilde{P}_{M})$.
 $\tilde{\rho}_{M}$ and $\tilde{P}_{M}$ are the energy density and pressure of matter.
@@
 In a flat FRW background the field equations are given by
-$3\tilde{H}^{2}=\kappa^{2}[\frac{1}{2}(\frac{d\tilde{\phi}}{d\tilde{t}})^{2}+V(\tilde{\phi})+\tilde{\rho}_{M}]$ (3.5)
-$\frac{d^{2}\tilde{\phi}}{d\tilde{t}^{2}}+3\tilde{H}\frac{d\tilde{\phi}}{d\tilde{t}}+\frac{\partial V(\tilde{\phi})}{\partial{\tilde{\phi}}}=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})$ (3.6)
-$\frac{d\tilde{\rho}_{M}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{M}+\tilde{P}_{M})=\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.7)
-where $\tilde{H}\equiv\frac{1}{\tilde{a}}\frac{d\tilde{a}}{\tilde{d}t}$. One can rewrite equation (3.6) as
-$\frac{d\tilde{\rho}_{\phi}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.8)
-where $\tilde{\rho}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})+V(\tilde{\phi}))$ is the energy density of the scalar field and $\tilde{P}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})-V(\tilde{\phi}))$ is the pressure of the scalar field Using the standard thermodynamic relation $\tilde{P}_{M}=\omega\tilde{\rho}_{M}$, one can easily find a relation between $\tilde{\rho}_{M}$ and $\tilde{\phi}$ from equation (3.7)
-$\tilde{\rho}_{M}=\tilde{\rho}_{M0}\tilde{a}^{-3(1+\omega)}\exp({\frac{\kappa}{\sqrt{6}}(1-3\omega)(\tilde{\phi}-\tilde{\phi}_{0})})$ (3.9)
+$3\tilde{H}^{2}=\kappa^{2}[\frac{1}{2}(\frac{d\tilde{\phi}}{d\tilde{t}})^{2}+V(\tilde{\phi})+\tilde{\rho}_{M}]$ (9)
+$\frac{d^{2}\tilde{\phi}}{d\tilde{t}^{2}}+3\tilde{H}\frac{d\tilde{\phi}}{d\tilde{t}}+\frac{\partial V(\tilde{\phi})}{\partial{\tilde{\phi}}}=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})$ (10)
+$\frac{d\tilde{\rho}_{M}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{M}+\tilde{P}_{M})=\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (11)
+where $\tilde{H}\equiv\frac{1}{\tilde{a}}\frac{d\tilde{a}}{\tilde{d}t}$. One can rewrite equation (10) as
+$\frac{d\tilde{\rho}_{\phi}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (12)
+where $\tilde{\rho}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})+V(\tilde{\phi}))$ is the energy density of the scalar field and $\tilde{P}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})-V(\tilde{\phi}))$ is the pressure of the scalar field Using the standard thermodynamic relation $\tilde{P}_{M}=\omega\tilde{\rho}_{M}$, one can easily find a relation between $\tilde{\rho}_{M}$ and $\tilde{\phi}$ from equation (11)
+$\tilde{\rho}_{M}=\tilde{\rho}_{M0}\tilde{a}^{-3(1+\omega)}\exp({\frac{\kappa}{\sqrt{6}}(1-3\omega)(\tilde{\phi}-\tilde{\phi}_{0})})$ (13)
 So the set of dynamical equations to be handled are
-$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}\bigg{]}$ (3.10)
-$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m})\tilde{\phi}^{\prime}$ (3.11)
-$\tilde{\rho}_{m}^{\prime}+3\tilde{\rho}_{m}=\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{m}\tilde{\phi}^{\prime}$ (3.12)
-$\tilde{\rho}_{r}^{\prime}+4\tilde{\rho}_{r}=0$ (3.13)
+$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}\bigg{]}$ (14)
+$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m})\tilde{\phi}^{\prime}$ (15)
+$\tilde{\rho}_{m}^{\prime}+3\tilde{\rho}_{m}=\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{m}\tilde{\phi}^{\prime}$ (16)
+$\tilde{\rho}_{r}^{\prime}+4\tilde{\rho}_{r}=0$ (17)
 where prime denotes the differentiation with respect to $N(=\ln{\tilde{a}})$.
-Note that since the radiation field is traceless the scalar field does not couple to radiation (3.4).
+Note that since the radiation field is traceless the scalar field does not couple to radiation (8).
 In order to solve these coupled nonlinear inhomogeneous equations, we introduce a set of dynamical variables [37, 38, 39]
-$x=\frac{\kappa\tilde{\phi}\prime}{\sqrt{6}},y=\frac{\kappa\sqrt{V}}{\sqrt{3}\tilde{H}},\lambda=-\frac{V_{,\tilde{\phi}}}{kV},~{}m=\frac{\kappa^{2}\tilde{\rho}_{m}}{3\tilde{H}^{2}},~{}r=\frac{\kappa^{2}\tilde{\rho}_{r}}{3\tilde{H}^{2}}$ (3.14)
-and rewrite the equations (3.10-3.13) in terms of these new variables forming a set of autonomous systems of first-order differential equations
+$x=\frac{\kappa\tilde{\phi}\prime}{\sqrt{6}},y=\frac{\kappa\sqrt{V}}{\sqrt{3}\tilde{H}},\lambda=-\frac{V_{,\tilde{\phi}}}{kV},~{}m=\frac{\kappa^{2}\tilde{\rho}_{m}}{3\tilde{H}^{2}},~{}r=\frac{\kappa^{2}\tilde{\rho}_{r}}{3\tilde{H}^{2}}$ (18)
+and rewrite the equations (14-17) in terms of these new variables forming a set of autonomous systems of first-order differential equations
 […]
-where, $\Gamma=\frac{VV_{\tilde{\phi}\tilde{\phi}}}{V_{\tilde{\phi}}^{2}}$. In terms of these new variables, equation (3.10) takes the form $1=x^{2}+y^{2}+m+r$.
+where, $\Gamma=\frac{VV_{\tilde{\phi}\tilde{\phi}}}{V_{\tilde{\phi}}^{2}}$. In terms of these new variables, equation (14) takes the form $1=x^{2}+y^{2}+m+r$.
 4 General class of $f(R)$ : $f(R)=R+\alpha_{n}R^{n}$
 In this work, we consider a very general class of modified gravity $f(R)=R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has a suitable dimension.
-The Starobinsky model of modified gravity corresponds to $n=2$ i.e., $f(R)=R+\alpha_{2}R^{2}$, where $\alpha_{2}\sim\frac{1}{M^{2}}$.
+The Starobinsky model of modified gravity corresponds to $n=2$ i.e., $f(R)=R+\alpha_{2}R^{2}$, where $\alpha_{2}\sim\frac{1}{M^{2}}$, where $M$ is in MeV scale.
 With this general form of $f(R)$, the corresponding potential is
-$V(\tilde{\phi})={\cal A}e^{-2\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\bigg{(}e^{\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}-1\bigg{)}^{\frac{n}{n-1}}$ (4.1)
+$V(\tilde{\phi})={\cal A}e^{-2\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\bigg{(}e^{\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}-1\bigg{)}^{\frac{n}{n-1}}$ (20)
 where ${\cal A}=\frac{(n-1)\alpha_{n}}{2\kappa^{2}}\big{[}\frac{1}{n\alpha_{n}}\big{]}^{\frac{n}{n-1}}$.
@@
 For $n\neq 2$, the potential increases sharply for negative $\tilde{\phi}$ and goes to 00 for positive $\tilde{\phi}$.
-In fig (1) we present the form of potentials for various values of $n$ with ${\cal A}=1$.
+In fig (1) we present the form of potentials for various values of $n$.
+The form of the potentials are independent of the choice of $\cal A$.
 Figure 1: Form of potentials for different $n$
@@
 For $n=2$, $f(R)$ takes the form $R+\alpha_{2}R^{2}$, which is the famous Starobinsky model. Substituting the value of $n$, the $\frac{d\lambda}{dN}$ equation becomes
-$\frac{d\lambda}{dN}=\sqrt{\frac{3}{2}}\lambda^{2}x\left[1-\frac{1}{\lambda}\sqrt{\frac{8}{3}}\right]$ (5.1)
-Note that the other equations in (3.15) remain the same.
-We solve the above equations numerically, with suitable initial conditions such that radiation-matter equality happens around redshift 3300 and matter in the present universe constitutes approximately 30% of the energy density.
-In Fig (2), we present different density parameters, scalar field $\tilde{\phi}$ and the equation of state $\omega_{\tilde{\phi}}=\frac{\tilde{P}_{\tilde{\phi}}}{\tilde{\rho}_{\tilde{\phi}}}$ with respect to $N(=\ln{\tilde{a}})$ The density parameters of different components mimic the behaviours in standard model of cosmology.
-From the plot it is evident that the energy density of $\tilde{\phi}$ field never dominates the evolution until recently.
-Figure 2: Evolution of density parameters, scalar field $\tilde{\phi}$ and $\omega_{\tilde{\phi}}$ in Starobinsky model
-$\tilde{\phi}$ remains positive along the entire evolution and does not vary much.
-$\omega_{\tilde{\phi}}$ remains constant for a long period and starts varying late, around $z\sim 7$ and finally approaches $-1$ at present.
+$\frac{d\lambda}{dN}=\sqrt{\frac{3}{2}}\lambda^{2}x\left[1-\frac{1}{\lambda}\sqrt{\frac{8}{3}}\right]$ (21)
+Note that the other equations in (19) remain the same.
+We solve the above equations numerically, with suitable initial conditions at the epoch of nucleosynthesis at $z\sim 10^{8}$.
+At such an early epoch radiation is the most dominating component compared to others .
+We have chosen $x=0.009,~{}y=2.2\times 10^{-15},~{}\lambda=-6\times 10^{-4}~{}{\text{and}}~{}r=0.9999$.
+Following these conditions radiation-matter equality happens around redshift 3300 and matter in the present universe constitutes approximately 30% of the energy density.
+In Fig (2), we present different cosmological parameters and the scalar field as they vary with redshift.
+The left panel presents different density parameters with respect to $\ln(1+z)$, The density parameters of matter and radiation follow their behavior of standard cosmology, while the density parameter of $\tilde{\phi}$ field never dominates the evolution until recently.
+To further understand the behavior of $\tilde{\phi}$, in the middle panel we plot $\omega_{\tilde{\phi}}(=\frac{\tilde{P}_{\tilde{\phi}}}{\tilde{\rho}_{\tilde{\phi}}})$ with respect to $\ln(1+z)$.
+$\omega_{\tilde{\phi}}$ remains constant for a long period and starts varying late and finally approaches $-1$ at present.
 So, in the late time the field behaves as dark energy.
 This behaviour of $\omega_{\tilde{\phi}}$ resembles the equation of state of the thawing quintessence field.
-We further investigate the potential and find that the scalar field rolls down the potential $V(\tilde{\phi})$ but does not reach the minima.
-This is due to the coupling between the scalar field and matter (see equations (3.11) and (3.12)).
-Plotting both the terms separately, we realised that the potential arising due to coupling with matter $V_{m}$ is far more dominant than the potential arising from modified gravity $V(\tilde{\phi})$.
-Hence the scalar field did not reach the minima.
-In Fig (3), we present the three different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$.
+However, since our main goal is to study the nature of the scalar field, in the right most panel we present $\tilde{\phi}$ with respect to $\ln(1+z)$.
+We observe that $\tilde{\phi}$ remains positive along the entire evolution and does not vary much.
+Figure 2: Evolution of density parameters, scalar field $\tilde{\phi}$ and $\omega_{\tilde{\phi}}$ in Starobinsky model(n=2).
+To understand the behaviour of the scalar field better, we investigate the potential $V(\tilde{\phi})$ and find that the scalar field rolls down the potential $V(\tilde{\phi})$ but does not reach the minima.
+This is due to the coupling between the scalar field and matter present on the right hand side of equation (10).
+Denoting $\frac{\partial V_{m}}{\partial\phi}=\kappa Q\tilde{\rho}_{m}$, one can write the potential $V_{m}$ due to the coupling term.
+Hence the scalar field rolls down the effective potential $V_{eff}(=V(\tilde{\phi})+V_{m}$).
+Plotting both the terms $V(\tilde{\phi})$ and $V_{m}$ separately, we realized that $V_{m}$ is far more dominant than $V(\tilde{\phi})$, the potential arising from modified gravity.
+Hence the scalar field rolls down the effective potential $V_{eff}(\approx V_{m})$ and did not reach the minima of $V(\tilde{\phi})$(see Fig (1)).
+In Fig (3), we present all the three different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$.
 Figure 3: Different potentials $V(\tilde{\phi})$, $V_{m}$ and $V_{eff}$ in Starobinsky model
-5.2 Case 2: $n\neq 2$, $f(R)=R+\alpha_{n}R^{n}$
+5.2 Case II: $n\neq 2$, $f(R)=R+\alpha_{n}R^{n}$
 Similarly, we have studied the other cases where $n$ takes positive values $3,4$ as well as negative values $-1,-2,-3,-4$ in the modified gravity form $f(R)=R+\alpha_{n}R^{n}$ and suitably changed the set of dynamical equations.
@@
 For $n=-2$, the $\frac{d\lambda}{dN}$ equation becomes
-$\frac{d\lambda}{dN}=\sqrt{\frac{3}{8}}x(4\sqrt{8/3}-6\lambda)(\sqrt{8/3}-\lambda)$ (5.2)
-along with the other equations in equation(3.15).
-Following similar methods as mentioned above, we solve the equations numerically, with suitable initial conditions.
-We did a similar analysis of the cosmological parameters as we have done in the case of $n=2$.
-Though the form of the potential of $V(\phi)$ in this case is different from the Starobinsky model,interestingly, we find that the behavior of the density parameters and the equation of state almost exactly matches the Starobinsky model over the entire period of evolution (Fig 4).
-This is because here also the potential $V_{m}$ arising due to matter coupling dominates, hence the effective potential $V_{eff}$ almost similarly matches the case of $n=2$ (Fig 5).
+$\frac{d\lambda}{dN}=\sqrt{\frac{3}{8}}x(4\sqrt{8/3}-6\lambda)(\sqrt{8/3}-\lambda)$ (22)
+along with the other equations in equation(19).
+Following similar methods as mentioned above, we solve the equations numerically, with suitable initial conditions at the epoch of nucleosynthesis.
+We set the initial conditions similar to the previous case ($x=0.008,~{}y=1.3\times 10^{-16},~{}\lambda=1.7~{}{\text{and}}~{}r=0.0.99997$ at $\ln(1+z)=18.8$.
+Note that negative initial values for $\lambda$ are not permissible in this case due to the form of potential.) We did a similar analysis of the cosmological parameters as we have done in the case of $n=2$.
+Though the form of the potential of $V(\phi)$ in this case is different from the Starobinsky model, interestingly, we find that the behavior of the density parameters and the equation of state are similar to the Starobinsky model over the entire period of evolution (Fig 4).
 The only difference is that the scalar field is negative for the entire period as it rolls down the potential.
+Further investigations on the modified gravity potential $V(\phi)$ and coupling potential $V_{m}$ reveal $V_{m}$ dominates over $V(\phi)$ like the previous case.
+Hence the scalar field rolls down the effective potential $V_{eff}(=V(\tilde{\phi})+V_{m})$ which is $\approx V_{m}$ (Fig 5).
+So, even though the form of the potential of $V(\phi)$ is different from the Starobinsky model, the cosmological parameters behaves similarly, but the scalar field remains negative.
 For all other values of n, ($\neq 2$), we get the same results.
@@
 The scalar field $\tilde{\phi}$ also couples with $\Lambda$ also, apart from matter.
-So, equations (3.10) and (3.11) changes to include $\Lambda$, while (3.12) and (3.13) remain unchanged.
+So, equations (14) and (15) changes to include $\Lambda$, while (16) and (17) remain unchanged.
 Also, we have one more coupled equation for $\Lambda$.
-$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}+\tilde{\rho}_{\Lambda}\bigg{]}$ (5.3)
-$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m}+4\tilde{\rho_{\Lambda}})\tilde{\phi}^{\prime}$ (5.4)
-$\tilde{\rho}_{\Lambda}^{\prime}=4\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{\Lambda}\tilde{\phi}^{\prime}$ (5.5)
+$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}+\tilde{\rho}_{\Lambda}\bigg{]}$ (23)
+$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m}+4\tilde{\rho_{\Lambda}})\tilde{\phi}^{\prime}$ (24)
+$\tilde{\rho}_{\Lambda}^{\prime}=4\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{\Lambda}\tilde{\phi}^{\prime}$ (25)
 To proceed with the analysis we define one more dynamical variable $l=\frac{\kappa^{2}\tilde{\rho}_{\Lambda}}{3\tilde{H}^{2}}$.
 In terms of all the dynamical variables, the new set of equations to be dealt with looks like
-$\frac{dx}{dN}$ $=\frac{x}{2}(3x^{2}+3-3y^{2}+r-3l)-3x+\sqrt{\frac{3}{2}}\lambda y^{2}$ (5.6)
+$\frac{dx}{dN}$ $=\frac{x}{2}(3x^{2}+3-3y^{2}+r-3l)-3x+\sqrt{\frac{3}{2}}\lambda y^{2}$ (26)
 $-\frac{1}{2}(1-x^{2}-y^{2}-r-l)-2l$
@@
 $\frac{dl}{dN}$ $=l[4x-(-3x^{2}+3y^{2}-r+3l-3)]$
-We perform a similar numerical analysis with the new set of equations for the potential given by (4.1) with different $n$ as has been done without $\Lambda$.
+We perform a similar numerical analysis with the new set of equations for the potential given by (20) with different $n$ as has been done without $\Lambda$.
 Surprisingly, we found that inclusion of the Cosmological constant does not effect the evolution of different components.
-And like the earlier scenario, the coupling term with matter $V_{m}$ is far more dominant than the potential $V(\tilde{\phi})$ arising out of modified gravity as well as the coupling with the cosmological constant $V_{l}$.
+Due to the inclusion of $\Lambda$, the scalar field couples with both matter and $\Lambda$ (equation (24)).
+So we have two coupling terms $V_{m}$ and $V_{l}$.
+But, even then, like the earlier scenario, $V_{m}$ is more dominant than both $V_{l}$ and $V(\tilde{\phi})$ of modified gravity.
 Hence driven mostly by the matter coupling term $V_{m}$, the evolution of different components remain similar.
@@
 Here we present only one case $n=2$ as an example.
-Figure 6: Evolution of different cosmological parameters and potentials including $\Lambda$ for $n=2$
+Figure 6: Evolution of different cosmological parameters and potentials including $\Lambda$ for $n=2$.
+The initial conditions in these plots are same as mentioned in section 5.1 along with $l=6\times 10^{-29}$ at $\log(1+z)=18.8$ for the new variable $l$.
 6 Results
-We have considered a very general class of modified gravity of the form $f(R)=R+\alpha_{n}R^{n}$ gravity, which is equivalent to Einstein gravity along with a scalar degree of freedom $\tilde{\phi}$ following the potential given by (4.1), when conformally transformed in Eintein frame.
-Depending on the behaviour of the potential we divided our cases in two categories $n=2$ (Case I) and $n\neq 2$ (Case 2).
-We investigated the evolution of different components like scalar field, radiation and matter in a background isotropic and homogeneous universe represented by FRW metric.
-After the whole analysis we found that in case $n=2$, we found that $\tilde{\phi}$ remains positive along the entire evolution and does not vary much.
+We have considered a very general class of modified gravity of the form $f(R)=R+\alpha_{n}R^{n}$ in an early universe around the epoch of nucleosynthesis.
+When conformally transformed in Eintein frame, modified gravity is equivalent to Einstein gravity along with a scalar degree of freedom $\tilde{\phi}$ following the potential given by (20).
+Depending on the behaviour of the potential we divided our cases in two categories $n=2$ (Case I) and $n\neq 2$ (Case II).
+We investigate the evolution of different components like scalar field, radiation and matter in a background isotropic and homogeneous universe represented by FRW metric.
+After the whole analysis we found that in the case of $n=2$, $\tilde{\phi}$ remains positive along the entire evolution and does not vary much.
 The matter - scalar field coupling does not allow the scalar field to reach the minima at zero.
-Such values of $\tilde{\phi}$ for the last two terms in action (2.4) provide an additional exponential suppression.
+Such values of $\tilde{\phi}$ for the last two terms in action (4) provide an additional exponential suppression.
 This makes the signature of the antisymmetric fields (both rank 2 and 3) heavily suppressed thereby failing to leave some imprints in the observable universe.
 The investigations for the case of $n\neq 2$ reveal $\tilde{\phi}$ to be negative all along (again matter-$\tilde{\phi}$ coupling not allowing $\tilde{\phi}$ to reach zero).
-Such values of $\tilde{\phi}$ would imply enhancement of the coupling for the last two terms in action (2.4) by an exponential factor.
+Such values of $\tilde{\phi}$ would imply enhancement of the coupling for the last two terms in action (4) by an exponential factor.
 This indicates that signatures of the antisymmetric fields of rank 2 and 3 should be present, which is contrary to the observation.
@@
 Thus, in the Einstein frame, $\tilde{\phi}$ couples with all the other fields through the conformal transformation of the metric.
-In this work, we first studied how different components like scalar field, radiation and matter evolve in a background isotropic and homogeneous universe represented by FRW metric.
-Further, we explore how the nature of the scalar field influences the coupling of the massless modes of higher rank antisymmetric fields present in the action (2.4).
+In this work we study in an early universe post reheating, how different components of energy evolve in a background isotropic and homogeneous universe represented by FRW metric.
+Further, we explore how the nature of the scalar field influences the coupling of the massless modes of higher rank antisymmetric fields present in the action (4).
 The coupling, in turn, is responsible for the observation of the possible imprints of massless modes of these antisymmetric fields.
 For this study, we have considered the form of $f(R)$ to be $R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has suitable dimensions.
-As shown in fig(1), the shape of the potential $V(\tilde{\phi})$ for this form is same for all values of $n$, apart from $n=2$.
-For $n\neq 2$, the potential is zero for positive $\tilde{\phi}$ and increases exponentially for $\tilde{\phi}<0$.
-The $n=2$ case, Starobinsky model, is different from the rest, as the potential increases and soon reaches a constant for positive $\tilde{\phi}$ but for negative $\tilde{\phi}$, the behaviour is the same as other cases.
+As shown in fig(1), the shape of the potential $V(\tilde{\phi})$ remain same for all values of $n\neq 2$ the potential is zero for positive $\tilde{\phi}$ and increases exponentially for $\tilde{\phi}<0$.
+But, the $n=2$ Starobinsky model is different from others, as the potential is positive for positive $\tilde{\phi}$ and negative for negative $\tilde{\phi}$ with a minima at $\tilde{\phi}=0$.
 The field equations with different kinds of matter and scalar field in the homogeneous and isotropic background are coupled nonlinear differential equations.
 We rewrite the equations as a set of autonomous system of first order differential equations to tackle the complexity.
-The autonomous system is solved numerically, with suitable initial conditions such that, following standard model of cosmology, the matter radiation equality occurs around redshift $z\sim 3300$ and the matter content is $30\%$ of the total in the present universe.
+The autonomous system is solved numerically, with suitable initial conditions at early epoch of nucleosynthesis.
 Though we studied different cases for $-4\geq n\geq 4$, we broadly divide them into two categories $n=2$ and $n\neq 2$, depending on the behaviour of the potential.
@@
 The effective potential exhibit similar nature for different values of $n$ and the variation of potential $V(\tilde{\phi})$ does not have any significant impact on the evolution of the parameters.
-Most interestingly, for different values of $n$, even though the different matter components, namely matter, radiation and scalar field evolve in the similar fashion with time, $\tilde{\phi}$ is positive only for $n=2$ and negative for $n\neq 2$, throughout.
+Most interestingly, for different values of $n$, even though the behaviour of the density parameters of different components, namely matter, radiation and scalar field remain similar, $\tilde{\phi}$ is positive only for $n=2$ and negative for $n\neq 2$, throughout.
 The value of $\tilde{\phi}$ depends on the slope of the potential.
-For $n=2$, $\tilde{\phi}$ remains positive as it rolls towards the minima ($\tilde{\phi}=0$).
+For $n=2$, $\tilde{\phi}$ remains positive as it rolls towards the minima ($\tilde{\phi}=0$) on the positive side of the potential.
 Where as, for $n\neq 2$, $\tilde{\phi}$ can roll only over negative value, as on the positive side the potential is zero.
 However, for both the cases $\tilde{\phi}$ can never reach zero due to coupling with matter.
-Hence even though both the categories agree with standard model of cosmology, only for Starobinsky model $(n=2)$, the last two terms of the action (2.4) get an extra exponential suppression.
+Hence even though both the categories agree with standard model of cosmology, only for Starobinsky model $(n=2)$, the last two terms of the action (4) get an extra exponential suppression.
 For $n\neq 2$, we obtain the enhancement of the coupling term instead of the suppression.